# Many-hypercube codes: High-rate quantum error-correcting codes for high-performance fault-tolerant quantum computation

### v1

@@ Appendix G Simulation of the bit-flip error model
To evaluate the performance of the decoders, we did the following numerical Monte Carlo simulation. We first prepare an error-free logical all-zero state of the C642 using the encoder in Fig. 5d. Then independent physical-qubit bit-flip errors are induced with probability $p_{\mathrm{flip}}$ per physical qubit. Finally, we ideally measure the physical qubits in the $Z$ basis and decode the outcomes. If all the logical-qubit outcomes are 0, the decoding succeeds, and otherwise it fails. We evaluated the logical failure probabilities by sufficiently many trials of this simulation.
Appendix H Error-detection gadgets
Figures 12a and 12b show the level-1 $Z$-error and $X$-error error-detection gadgets, respectively, used in the level-2 zero-state encoder in Fig. 3b. Figures 12c and 12d show the level-2 $Z$-error and $X$-error error-detection gadgets, respectively, used in the level-3 zero-state encoder.
Appendix I Simulation of logical CNOT gates
To evaluate the performance of full transversal logical CNOT gates, we did the numerical Monte Carlo simulation shown in Fig. 13. From many trials of this simulation, we estimate the error probability and its standard error of 10 sets of full transversal logical CNOT gates, which are denoted by $p_{10}$ and $\Delta_{10}$, respectively. Then we evaluate those for one set, which are denoted by $p_{1}$ and $\Delta_{1}$, as $p_{1}=1-(1-p_{10})^{1/10}$ and $\Delta_{1}=\frac{\Delta_{10}}{10}(1-p_{10})^{1/10-1}$. Since each set includes $K^{(L)}$ logical CNOT gates at level $L$, we finally evaluate the error probability and its standard error per logical CNOT gate denoted by $p_{\mathrm{CNOT}}$ and $\Delta_{\mathrm{CNOT}}$ as $p_{\mathrm{CNOT}}=1-(1-p_{1})^{1/K^{(L)}}$ and $\Delta_{\mathrm{CNOT}}=\Delta_{1}/K^{(L)}(1-p_{1})^{1/K^{(L)}-1}$.
Appendix J Logical phase and non-Clifford gates
The full transversal logical phase and non-Clifford gates, $S$ and ${R_{Y}(\pi/4)}$, are performed with logical $|Y\rangle=(|0\rangle+i|1\rangle)/\sqrt{2}$ and $|H\rangle=\cos(\pi/8)|0\rangle+\sin(\pi/8)|0\rangle$, respectively, as shown in Figs. 14a and  14b. For the non-Clifford gate, we need partial transversal logical Hadamard gates depending on the measurement outcomes. We can also perform partial transversal logical phase and non-Clifford gates by replacing the full transversal logical Hadamard gates with partial ones, because then the two CNOT gates on the qubits without Hadamard gates are cancelled out.
Appendix K Partial transversal logical Hadamard gates
The partial transversal logical Hadamard gates are performed with a logical partial |+ket|$\rangle$| + ⟩ state and full transversal logical CNOT and Hadamard gates, as shown in Fig. 15a. The logical partial |+ket|$\rangle$| + ⟩ states are prepared fault-tolerantly by the arbitrary-state encoder in Fig. 5e and the standard state verification shown in Fig. 15b.
Appendix L Logical SWAP gates
The level-1 encoded SWAP gates are performed by physical SWAP gates as shown in Fig. 5c. At higher levels, e.g., level 3, simultaneous physical SWAP gates $\{\mathrm{Q}_{i^{\prime},j^{\prime},1}\leftrightarrow\mathrm{Q}_{i^{\prime},j^{\prime},3}|i^{\prime}=1,\ldots,6,j^{\prime}=1,\ldots,6\}$ result in simultaneous logical SWAP gates $\{\mathrm{Q}_{i,j,1}^{(3)}\leftrightarrow\mathrm{Q}_{i,j,2}^{(3)}|i=1,\ldots,4,j=1,\ldots,4\}$. Similarly, the other simultaneous logical SWAP gates, such as $\{\mathrm{Q}_{i,3,k}^{(3)}\leftrightarrow\mathrm{Q}_{i,4,k}^{(3)}|i=1,\ldots,4,k=1,\ldots,4\}$ and $\{(\mathrm{Q}_{1,j,k}^{(3)},\mathrm{Q}_{2,j,k}^{(3)})\leftrightarrow(\mathrm{Q}_{3,j,k}^{(3)},\mathrm{Q}_{4,j,k}^{(3)})|j=1,\ldots,4,k=1,\ldots,4\}$, are performed by corresponding simultaneous physical SWAP gates, such as $\{\mathrm{Q}_{i^{\prime},4,k^{\prime}}\leftrightarrow\mathrm{Q}_{i^{\prime},6,k^{\prime}}|i^{\prime}=1,\ldots,6,k^{\prime}=1,\ldots,6\}$ and $\{(\mathrm{Q}_{1,j^{\prime},k^{\prime}},\mathrm{Q}_{2,j^{\prime},k^{\prime}},\mathrm{Q}_{3,j^{\prime},k^{\prime}})\leftrightarrow(\mathrm{Q}_{4,j^{\prime},k^{\prime}},\mathrm{Q}_{5,j^{\prime},k^{\prime}},\mathrm{Q}_{6,j^{\prime},k^{\prime}})|j^{\prime}=1,\ldots,6,k^{\prime}=1,\ldots,6\}$. In general, we can perform simultaneous logical SWAP gates between the logical qubits corresponding to hypercubes on two parallel hyperplanes by simultaneous physical SWAP gates.
Appendix M Logical CNOT gates
Partial transversal logical CNOT gates between two registers (inter-block logical CNOT gates) can be performed using two ancilla registers in the same partial |+ket|$\rangle$| + ⟩ state and full transversal logical CNOT and Hadamard gates, as shown in Fig. 16a. Using logical SWAP gates, we can perform a logical CNOT gate on an arbitrary logical-qubit pair between the two registers. On the other hand, logical CNOT gates in a register (intra-block logical CNOT gates) can be performed with an additional ancilla register in the logical all-zero state, logical SWAP gates, and full and partial transversal logical CNOT gates, as shown in Fig. 16b.

### v3

@@ Appendix E Simulation of the bit-flip error model
To evaluate the performance of the decoders, we did the following numerical Monte Carlo simulation. We first prepare error-free logical zero states of the many-hypercube code using the encoder in Fig. 1A. Then independent physical-qubit bit-flip errors are induced with probability $p_{\mathrm{flip}}$ per physical qubit. Finally, we ideally measure the physical qubits in the $Z$ basis. This quantum circuit includes measurements only at the end. Therefore, we can use the fast sampler of Stim [67] to obtain the physical-qubit measurement outcomes. Then we decode the outcomes using our homemade python codes implementing the above-mentioned decoding methods. If all the logical-qubit outcomes are 0, the decoding succeeds, and otherwise fails. We evaluated the decoding error probabilities by sufficiently many trials of this simulation.
Appendix F Simulation of logical CNOT gates
To evaluate the performance of full transversal logical CNOT gates in the circuit-level noise model, we did the numerical Monte Carlo simulation shown in Fig. 10. In this simulation, we first prepare two error-free logical Bell-state blocks using the zero-state encoder in Fig. 1A and error-free physical CNOT and Hadamard gates. Then, we perform faulty full transversal logical CNOT gates implemented by faulty full transversal physical CNOT gates on the first code blocks of the two Bell-state blocks. This is followed by ECT gadgets in which physical operations are faulty according to the noise model. The faulty logical CNOT gates with faulty ECT gadgets are repeated 10 times. After that, the Bell states are disentangled and converted to the logical zero states by error-free physical CNOT and Hadamard gates. Finally, the logical zero states are ideally measured in the $Z$ basis and the measurement outcomes are decoded by the proposed minimum-distance decoder. Unlike the above simulation of bit-flip errors, this simulation requires mid-circuit measurements and feed-forward operations based on the measurement outcomes. Therefore we use the slow TableauSimulator of Stim [67] for stabilizer quantum-circuit simulation parts. From many trials of this simulation, we estimate the error probability and its standard error of 10 sets of full transversal logical CNOT gates, which are denoted by $p_{10}$ and $\Delta_{10}$, respectively. Then we evaluate those for one set, which are denoted by $p_{1}$ and $\Delta_{1}$, as $p_{1}=1-(1-p_{10})^{1/10}$ and $\Delta_{1}=\frac{\Delta_{10}}{10}(1-p_{10})^{1/10-1}$. Since each set includes $K^{(L)}$ logical CNOT gates at level $L$, we finally evaluate the error probability and its standard error per logical CNOT gate denoted by $p_{\mathrm{CNOT}}$ and $\Delta_{\mathrm{CNOT}}$ as $p_{\mathrm{CNOT}}=1-(1-p_{1})^{1/K^{(L)}}$ and $\Delta_{\mathrm{CNOT}}=\Delta_{1}/K^{(L)}(1-p_{1})^{1/K^{(L)}-1}$.
Figure 10: Simulation of logical CNOT gates. We first prepare two sets of error-free logical Bell states encoded with the level-$L$ many-hypercube code in four registers. Then we perform 10 times faulty full transversal logical CNOT gates (implemented by faulty full transversal physical CNOT gates) followed by faulty ECT gadgets on the first and third registers in the circuit-level noise model. Finally, we disentangle the logical states by error-free operations and ideally measure all the physical qubits in the $Z$ basis. The measurement outcomes are decoded by the minimum distance decoder. If all the logical outputs are 0, the 10 sets of full transversal logical CNOT gates succeed, and otherwise fail.
Appendix G Results at level 2 using the Steane method for error detection
The crosses in Fig. 11 show the results at level 2 using the Steane method in error-detection gadgets instead of the flag-based method. As found from Fig. 11, the Steane method leads to no performance improvement and a larger space overhead, compared to the flag-based method. This is the reason why we use the flag-based method at level 2.
Figure 11: Results at level 4 with the zero-state encoder in Fig. 5B. (A) Space overheads for the zero-state encoder. (B) Logical CNOT error probability.
Appendix H Results at level 4 using the encoder in Fig. 5B
In this work, we propose the level-4 zero-state encoder in Fig. 5G to improve the performance compared to the case of the encoder in Fig. 5B. Here we present the results in the latter case. Figure 12A shows the space overheads at level 4. From this, the encoder in Fig. 5B with the most stringent condition for error detection, $L_{D}=1$ (see Appendix D), leads to the rapid increase of the space overhead with respect to $p_{\mathrm{circ}}$ (shown by squares in Fig. 12A) due to low acceptance probabilities. Relaxing the condition to $L_{D}=2$, the overheads (shown by triangles in Fig. 12A) become similar to those with the encoder in Fig. 5G (shown in circles in Fig. 12A and also in Fig. 6B). However, then the logical CNOT error probability (shown by triangles in Fig. 12B) is much worse than that with the encoder in Fig. 5G (shown in circles in Fig. 12B and also in Fig. 6A). This is the reason why we propose the encoder in Fig. 5G at level 4.
Figure 12: Results at level 4 with the zero-state encoder in Fig. 5B. (A) Space overheads for the zero-state encoder. (B) Logical CNOT error probability.
Appendix I Details and flowcharts of the proposed level-by-level minimum distance decoding
Here we explain our proposed decoding method in more detail using flowcharts. As an example, we focus on the level-3 case.
Figure 13 shows the whole process of the decoding in the level-3 case. Since the subroutines Sub1 and Sub2 are well explained in Appendix D, here we explain the subroutine Sub3 in detail.
Figure 14 shows the process of Sub3. The subroutine Sub31 for $b_{2}=6$ is shown in Fig. 15, where $N_{1}$ to $N_{5}$ denote the numbers of the minimum-distance candidates of the first to fifth level-2 blocks, respectively, selected in Sub2. If the total number of candidates, $N_{1}N_{2}N_{3}N_{4}N_{5}$, is larger than a preset threshold $N_{\mathrm{th}3}$, we reduce the number by randomly choosing one of the candidates from the level-2 block with the largest $N$, in order to keep the decoding time short. In this work, we set $N_{\mathrm{th}3}$ to $10^{5}$. Then we determine the encoded bit string of the sixth level-2 block using the $n_{1}$-th to $n_{5}$-th minimum-distance candidates of the first to fifth level-2 blocks, respectively, according to the parity-check condition ($Z$-stabilizer condition). More concretely, the encoded bit string of the sixth level-2 block is determined as follows:
$x_{i^{\prime},j^{\prime},6}^{(2)}=x_{i^{\prime},j^{\prime},1}^{(2)}+x_{i^{\prime},j^{\prime},2}^{(2)}+x_{i^{\prime},j^{\prime},3}^{(2)}+x_{i^{\prime},j^{\prime},4}^{(2)}+x_{i^{\prime},j^{\prime},5}^{(2)},$ (23)
In general, the level-2 encoded bit string determined above is not included in the minimum-distance candidates of this block selected in Sub2. Therefore, in the subroutine Sub32, we evaluate the distance of this encoded bit string. Figure 16 shows Sub32 for $b_{2}=6$. In Fig. 16, $M_{1}$ to $M_{6}$ denote the numbers of the minimum-distance candidates of the first to sixth level-1 blocks, respectively, selected in Sub1 in the sixth level-2 block. If $M_{1}+M_{2}+M_{3}+M_{4}+M_{5}+M_{6}$ is larger than a preset threshold $M_{\mathrm{th}2}$, we reduce the number by randomly choosing one of the candidates from the level-1 block with the largest $M$. In this work, we set $M_{\mathrm{th}2}$ to 6. We first determine the encoded bit strings of the five level-1 blocks other than the $b$-th level-1 block in the sixth level-2 block using the $m_{b}$-th minimum-distance candidate of the $b$-th level-1 block selected in Sub1 and the level-2 encoding bit string determined in Sub31 according to the parity-check condition and the definition of the encoded $Z$ operator. More concretely, when $b=1$, the encoded bit strings of the five level-1 blocks are determined as follows:
$x_{i^{\prime},2,6}^{(1)}$ $=x_{i^{\prime},1,6}^{(1)}+x_{i^{\prime},1,6}^{(2)},$ (24)
$x_{i^{\prime},3,6}^{(1)}$ $=x_{i^{\prime},2,6}^{(1)}+x_{i^{\prime},2,6}^{(2)},$ (25)
$x_{i^{\prime},5,6}^{(1)}$ $=x_{i^{\prime},2,6}^{(1)}+x_{i^{\prime},1,6}^{(2)}+x_{i^{\prime},2,6}^{(2)}+x_{i^{\prime},3,6}^{(2)}+x_{i^{\prime},4,6}^{(2)},$ (26)
$x_{i^{\prime},4,6}^{(1)}$ $=x_{i^{\prime},5,6}^{(1)}+x_{i^{\prime},3,6}^{(2)},$ (27)
$x_{i^{\prime},6,6}^{(1)}$ $=x_{i^{\prime},5,6}^{(1)}+x_{i^{\prime},4,6}^{(2)},$ (28)
In general, the level-1 encoded bit strings determined above are not included in the minimum-distance candidates of the level-1 blocks selected in Sub1. Therefore, we evaluate the distances of the encoded bit strings. This distance evaluation can easily be achieved at level 1. By summing the distances of the six level-1 blocks, we obtain the distance of the sixth level-2 block. In Sub32, we finally select the minimum distance of the sixth level-2 block among all the choices of a level-1 block and its minimum-distance candidates. As shown in Fig. 15, then we evaluate the distances of the level-3 codewords by summing the distances of their six level-2 blocks. We finally select level-3 encoded bit strings with minimum distance, as shown in Fig. 14.
The decoding of the level-4 many-hypercube code is done in a similar manner to the above level-3 case. In the level-4 case, we set the level-4 threshold $N_{\mathrm{th}4}$ corresponding to the above level-3 threshold $N_{\mathrm{th}3}$ to the same value $10^{5}$ and the level-3 threshold $M_{\mathrm{th}3}$ corresponding to the above level-2 threshold $M_{\mathrm{th}2}$ to 12.